OpenStudy (anonymous):

write the decimal as a fraction or a mixed number in simplest form 3.4

OpenStudy (anonymous):

write it "over 1" and then multiply both the numerator and denominator to get whole numbers on the top and bottom, here we can multiply by 10 \[3.4 = \frac{3.4}{1} = \frac{34}{10}\] \[\frac{34}{10} =\frac{17}{5}\]

OpenStudy (radar):

3 2/5 or 17/5

OpenStudy (anonymous):

the last step is to simplify the fraction by dividing until they have no common factors

OpenStudy (radar):

3 4/10 = 3 2/5=17/5